I need to determine the domain of the given function
f(x) = 4x divided by x^ -x -12

Answer by lwsshak3(11628)

I need to determine the domain of the given function
f(x) = 4x divided by x^ -x -12
...
f(x)=4x/(x^2-x-12)
To determine domain, set denominator = 0, then solve for x
x^2-x-12=0
(x-4)(x+3)=0
x-4=0
or
x=-3
Domain:
x≠4 or x≠-3
in interval notation.
(-∞,-3) U (-3,4) U (4,∞)